# PromptIntern: Saving Inference Costs by Internalizing Recurrent Prompt during Large Language Model Fine-tuning

### v1

@@ 5.3 Ablation Study
  • PromptIntern w/o Example Absorption, in which we omit the example absorption for retrieving and internalizing few-shot examples during fine-tuning.
@@
The overall results is shown in Table 5. When comparing PromptIntern with PromptIntern w/$\tau_{tmp}=0.3$, we observe an average of 2.4% drop on performance but a 3.7x compression on tokens across all three datasets. This highlights the balance between compression rate and accuracy performance. When comparing our with our w/o Example Absorption, we observe a significant performance drop in the latter variant, despite both approaches utilizing the same number of tokens for inference. This outcome highlights the importance of example absorption in internalizing essential information during the fine-tuning process. When comparing PromptIntern with PromptIntern w/o Template Compression, we note that adding the template compression saves an average of 280 tokens across the datasets but experiences an average performance drop of 5%. This demonstrates that while totally internalizing the template into model parameters significantly reduces token usage, it requires a trade-off in terms of inference performance.
@@ 5.5 Analysis on Examples Retrieval Bank
Table 6 examines the impact of varying proportion of the training set used for constructing relevant examples in the examples retrieval bank. The options tested include using 25%, 50%, and 100% of the training set. The results clearly show a trend where increasing the percentage of the training set used in the examples retrieval bank correlates with improved performance. This suggests that larger examples retrieval bank provides a richer set of few-shots for the model to learn from, thereby enhancing its ability to generalize and perform accurately across tasks.
@@ 6 Conclusion
In this paper, we propose PromptIntern, a novel method for prompt internalization that internalizes repetitive prompt knowledge into LLMs parameters. We develop specific compression strategies for different components of the prompt, accompanied by a tailored progressive fine-tuning pipeline. Extensive experiments demonstrate that our method maintains comparable performance effectiveness while significantly accelerating inference speed and reducing token usage.
@@ 7 Limitations
While PromptIntern significantly reduces costs during the inference stage, the progressive fine-tuning approach incurs additional computational expenses during training. Specifically, our methodology demands substantial manual intervention for pre-processing and parameter adjustments throughout the fine-tuning process. Moreover, the current evaluation of our method is limited to a single task, specifically NL2Code. This restricts our understanding of its generalizability and effectiveness across a broader range of tasks. In future work, we plan to conduct extensive evaluations on more complex and varied tasks, such as long document summarization and question answering within specialized technical domains.
@@ 8 Ethics Statement
This research does not raise any ethical concerns. We obtained data only from publicly available sources where users have consented to the public sharing of their posts. We have conducted a thorough assessment to ensure that our research does not pose any potential harm.
@@ A.2 Implementation Details
Baseline Settings
Parameter Settings for PromptIntern
2) Number of $\tau_{tmp}$ for template compression: For the prompt compression baseline experiments, we set the final template rate to 0.3, which is used in the last stage of fine-tuning as well as inference. For the other experiments and ablation studies, we set the final template rate to 0 to achieve fully internalization.
Cost Evaluation
Computational Resource
@@ A.3 Comparison with Gist Tokens
We report the comparison result of PromptIntern with Gist Tokens Mu et al. (2024) on Table 7. Gist Tokens showcases consistent performance, with notable results in NL2Bash where it achieves a BLEU score of 22.7, suggesting a moderate alignment with the dataset’s requirements. In contrast, PromptIntern demonstrates superior performance across all metrics and datasets, particularly excelling in the NL2Bash dataset with a BLEU score of 31.6 and maintaining similar efficiency in token usage. The results demonstrate the our approach significantly outperforms the Gist token while conducting overall the same compression rate.
A.4 Experiments on Inference Speed
The experimental results presented in Table 8 illustrate the low latency characteristics of PromptIntern during inference across three datasets, MBPP,NL2F, and NL2Bash. Specifically, for the MBPP dataset, PromptIntern achieves an inference speed of 4.17 s/instance on the GPT-4 model, closely aligning with the 4.36 s/instance observed in the no template setup and far surpassing the more resource-intensive template with 5-shots configuration at 10.21 s/instance. In the NL2F dataset, PromptIntern similarly demonstrates its efficiency with an inference speed of 2.15 s/instance for GPT-4, which is nearly equivalent to the 2.12 s/instance observed without any template and significantly outperforms the elaborate template with 10-shots configuration, which achieves 12.47 s/instance. The experimental results outlined in the table also highlight the efficiency of PromptIntern in the NL2Bash dataset. Notably, for GPT-4 under the NL2Bash benchmark, PromptIntern maintains a competitive inference speed of 1.18 s/instance, matching the performance seen in the no template scenario and markedly better than the template with 5-shots setup, which records a slower speed of 4.46 s/instance. The result across three NL2Code benchmarks highlight PromptIntern ’s capability to maintain competitive inference speeds while minimizing latency efficiently.
Appendix B Example Demonstration
We demonstrate an example on how we schedule and pre-process an input prompt through both template compression and example absorption in Figure 3.
During the initial phase of fine-tuning, the input prompt will fully incorporate both the template and 10-shot examples for the NL2F dataset. After a specified number of training iterations, the template will undergo a compression at a rate of 0.3, and the number of examples will be reduced to five. This modified prompt is then used for the intermediate stage of fine-tuning. In the final phase, the template and few-shot examples are completely removed from the training prompt. It is important to note that the query remains unchanged throughout the entire progressive fine-tuning process. The prompt used in the last stage, which consists solely of the query, will also serve as the input for subsequent model inference. This method enables the fine-tuned language model to perform zero-shot inference without the need for a instruction or document template.
@@ Appendix B Example Demonstration
Appendix C Prompts
C.1 GPT-4 Generation (Baseline) instruction
@@
C.2 Prompts of PromptIntern on NL2Code
Please refer to Figures 5,6,7 for detailed prompts.

### v2

@@ 5.4 Ablation Study
  • PromptIntern w/o Example Absorption, where we omit the example absorption for retrieving and internalizing few-shot examples during fine-tuning.
@@
The overall results is shown in Table 3. When comparing PromptIntern with PromptIntern w/$\tau_{tmp}=0.3$, we observe an average of 2.4% drop on performance but a 3.7x compression ratio on token usages across all three datasets. This highlights the balance between compression rate and accuracy performance. When comparing our with our w/o Example Absorption, we observe a large performance drop in the latter variant, despite both approaches utilizing the same number of tokens for inference. The result demonstrates the importance of example absorption in internalizing essential information during the fine-tuning progress. In addition, when comparing PromptIntern with PromptIntern w/o Template Compression, we note that adding the template compression saves an average of 280 tokens across the datasets but experiences an average performance drop of 5%. The result above demonstrates that while internalizing the template into model parameters reduces token usage, it requires a trade-off in terms of inference performance.
@@ 5.6 Analysis on Examples Retrieval Bank
Table 4 examines the impact of varying proportion of the training set used for constructing relevant examples in the examples retrieval bank. The options tested include using 25%, 50%, and 100% of the training set. The results show a trend where increasing the percentage of the training set used in the examples retrieval bank correlates with improved performance. This suggests that larger examples retrieval bank provides a richer set of few-shots for the model to learn from, thereby enhancing its ability to perform accurately across tasks.
Additional Experiments.
In Appendix A.4, we explain why performing direct inference with a compressed prompt, while using the full prompt during training, is less effective compared to the progressive fine-tuning strategy employed by PromptIntern. Also, to provide a comprehensive analysis of PromptIntern’s efficiency, we evaluate the inference speed and calculate the overall monetary cost during inference (details in Appendix A.5). In addition, to show that the principle of PromptIntern can generalize beyond the NL2Code domain, we evaluate PromptIntern on an additional task, which is presented in Appendix A.6.
6 Discussion
Table 5: Time Overhead of PromptIntern during fine-tuning.
Model MBPP NL2F NL2Bash
PromptIntern Template with 5-shots PromptIntern Template with 10-shots PromptIntern Template with 10-shots
GPT-4 01h 38m 54s 02h 13m 07s 03h 46m 15s 04h 32m 12s 03h 23m 18s 04h 17m 28s
GPT-3.5 01h 19m 03s 01h 48m 55s 03h 02m 29s 03h 44m 46s 02h 23m 19s 03h 09m 16s
Llama2-13B 00h 45m 27s 01h 14m 10s 01h 36m 17s 02h 16m 21s 01h 12m 48s 01h 48m 03s
6.1 Training Overhead of PromptIntern
As PromptIntern is specifically designed to enhance the efficiency of LLMs during inference, the multi-stage design of progressive fine-tuning may introduce additional computational overhead. It is crucial to manage this overhead during the training phase to ensure that PromptIntern remains scalable and applicable in real-world, large-scale scenarios.
To illustrate the overhead of PromptIntern, we provide a detailed breakdown of the time overhead incurred during the training phase. We compare our method to the baseline template with k-shots, which represents common use cases. Our evaluation includes the complete end-to-end process, covering dataset import and training times. All experiments were conducted on an A100x1-80G GPU. As shown in Table 5, PromptIntern consistently requires less time for both data preparation and training compared to the baseline. This efficiency is primarily due to the reduced number of input tokens per training instance during progressive fine-tuning. Furthermore, as outlined in our experimental settings, the number of training epochs for PromptIntern matches that of the direct fine-tuning baselines, ensuring no additional computational cost from extra training steps.
6.2 General model ability
Although progressive fine-tuning ensures high accuracy for domain-specific tasks, concerns may arise regarding whether this multi-step fine-tuning approach could negatively impact the overall model’s generalizability. Several studies have examined fine-tuning for LLMs. Wei et al. (2021) shows that multi-task fine-tuning enhances zero-shot and ICL capabilities. Mosbach et al. (2023) finds that few-shot fine-tuning preserves out-of-domain generalization similar to ICL settings. However, Wang et al. (2022) reveals that fine-tuning may overly adapt models to task-specific formats, reducing flexibility for new tasks. Luo et al. (2023a) also explore catastrophic forgetting during continual fine-tuning.
In this work, PromptIntern is designed to enhance the efficiency of fine-tuned LLMs during inference while maintaining task-specific performance. Our approach is particularly suited for scenarios where LLMs are fine-tuned and deployed for domain-specific tasks. Although PromptIntern is currently limited to fine-tuning a single task, recent inference optimization techniques Ye et al. (2023); Wang et al. (2023); Sheng et al. (2023) enable concurrent fine-tuning of multiple LoRA adapters while sharing a single backbone model. This allows for efficient adaptation to multiple tasks, reducing the number of parameters requiring fine-tuning. We will leave the integration of PromptIntern with these approaches in our future work.
We also provide a detailed discussion on preventing model overfitting in Appendix B.1.
@@ 7 Conclusion
In this paper, we propose PromptIntern, a prompt internalization method that internalizes repetitive prompt knowledge into LLMs parameters. We develop specific compression strategies for different components of the prompt, accompanied by a tailored progressive fine-tuning pipeline. Extensive experiments demonstrate that our method maintains comparable performance effectiveness while accelerating inference speed with less token usage.
@@ 8 Limitations
Through extensive experiments in this paper, PromptIntern has demonstrated a strong ability to reduce model costs during inference. However, as shown in Table 2, PromptIntern still exhibits a performance gap in accuracy compared to the use of original prompts (i.e., template with k-shots). Also, while we empirically validate the effectiveness of PromptIntern, a theoretical analysis of model parameter updates and the training pipeline complexity is still required. In addition, although the principle of PromptIntern can be generalized to most downstream NLP tasks (as demonstrated in Appendix A.6, further empirical verification is needed on more advanced tasks. In future work, we plan to conduct more evaluations on several complex tasks, including long-document summarization, question-answering in specialized technical domains, etc.
@@ 9 Ethics Statement
All datasets used in this paper are publicly available and have been reviewed to ensure they do not contain any personally identifiable information or offensive content. Additionally, experiments were conducted on computational clusters with NVIDIA A100 GPUs. It is important to note that this could have an environmental impact, and the carbon footprints were monitored in real time.
@@ A.2 Implementation Details
Baseline Settings
Parameter Settings for PromptIntern
2) Number of $\tau_{tmp}$ for template compression: For the prompt compression baseline experiments, we set the final template rate to 0.3, which is used in the last stage of fine-tuning as well as inference. For the other experiments and ablation studies, we set the final template rate to 0 to achieve full internalization.
@@ A.3 Comparison with Gist Tokens
We report the comparison result of PromptIntern with Gist Tokens Mu et al. (2024) on Table 6. Gist Tokens showcases consistent performance, with notable results in NL2Bash where it achieves a BLEU score of 22.7, suggesting a moderate alignment with the dataset’s requirements. In contrast, PromptIntern demonstrates superior performance across all metrics and datasets, particularly excelling in the NL2Bash dataset with a BLEU score of 31.6 and maintaining similar efficiency in token usage. The results demonstrate that our approach significantly outperforms the Gist token while conducting overall the same compression rate.
A.4 Effectiveness of progressive fine-tuning in PromptIntern
In this experiment, we compare PromptIntern with the method of direct fine-tuning with a full-loaded prompt (template plus few-shot examples) followed by inferencing with queries only (designated as Template with k-shots* in table 8). We use the comparison to demonstrate the effectiveness of progressive fine-tuning for updating model parameters properly. The result is shown in Table 8. We can clearly observe that
Table 8: Demonstration of progressive fine-tuning in PromptIntern
Dataset (Inference on GPT3.5) MBPP NL2F NL2Bash
Template with k-shots 69.1 71.7 63.2
PromptIntern 76.6 78.4 67.7
Template with k-shots* has a large performance degradation compared to PromptIntern. This indicates that fine-tuned LLMs struggle to establish a proper connection between the full-length prompts used in training and the query-only prompts used during inference. It also motivates the development of the progressive fine-tuning strategy in PromptIntern. Beyond empirical experiments demonstration, we will leave the theoretical proof of the effectiveness of PromptIntern (Algorithm 1) in our future work.
@@ A.4 Effectiveness of progressive fine-tuning in PromptIntern
A.5 Experiments on Inference Speed
The experimental results presented in Table 7 illustrate the low latency characteristics of PromptIntern during inference across three datasets, MBPP, NL2F, and NL2Bash. Specifically, for the MBPP dataset, PromptIntern achieves an inference speed of 4.17 s/instance on the GPT-4 model, closely aligning with the 4.36 s/instance observed in the no template setup and far surpassing the more resource-intensive template with 5-shots configuration at 10.21 s/instance. In the NL2F dataset, PromptIntern similarly demonstrates its efficiency with an inference speed of 2.15 s/instance for GPT-4, which is nearly equivalent to the 2.12 s/instance observed without any template and significantly outperforms the elaborate template with 10-shots configuration, which achieves 12.47 s/instance. The experimental results outlined in the table also highlight the efficiency of PromptIntern in the NL2Bash dataset. Notably, for GPT-4 under the NL2Bash benchmark, PromptIntern maintains a competitive inference speed of 1.18 s/instance, matching the performance seen in the no template scenario and markedly better than the template with 5-shots setup, which records a slower speed of 4.46 s/instance. The result across three NL2Code benchmarks highlights PromptIntern ’s capability to maintain competitive inference speeds while minimizing latency efficiently.
A.6 Evaluating PromptIntern on GSM8K
To demonstrate the generalization ability of PromptIntern on other domain tasks, we also test on the GSM8K dataset Cobbe et al. (2021). We use the same experiment settings stated in our experiment setups to compare with the prompt compression baselines. The results are shown in Table 9. The result demonstrates that, under the same compression rate for inference, PromptIntern outperforms other compression baselines by 3%-34% for the arithmetic reasoning task.
Table 9: Comparison of prompt compression baselines on GSM8K.
Methods (Inference on GPT 3.5) E.M. Tokens
Selective Context 63.5 443
LLMLingua 83.2 452
LLMLingua-2 81.9 458
PromptIntern 85.7 458
Appendix B Additional Discussions
B.1 Prevention of Model Over-fitting
To prevent LLMs from over-fitting due to lengthy input prompts, such as long templates, and to mitigate over-fitting during multi-stage fine-tuning, we have implemented several strategies:
  • Prompt Selection: For each dataset, we utilize default prompts provided by the authors or sourced from widely recognized papers. In our experiments, these prompts are applied for the baseline template with 5/10 shots, which outperforms other approaches and aligns with results from sources like Papers With Code. This ensures that our prompts will not cause performance degradation due to over-fitting input.
  • Model Checkpointing: To mitigate over-fitting during fine-tuning, particularly due to excessively lengthy training epochs, we implemented a model checkpointing strategy. We save the model state every 200 training steps and evaluate each checkpoint on a separate validation dataset. This allows us to track performance changes throughout the whole training process. By comparing checkpoints, we identify the optimal iteration that achieves the best results, determining the appropriate number of training steps and epochs for our experiments.
  • Validation Monitoring: Training is halted using an early stopping technique when the validation loss begins to rise or ceases to decline, indicating potential over-fitting. Additionally, we manually monitor the training and validation losses against the number of training steps for each experiment of PromptIntern. This visualization helps ensure that each model avoids training over-fitting by providing a clear depiction of the training dynamics and enabling timely adjustments.
Appendix C Example Demonstration
We demonstrate an example of how we schedule and pre-process an input prompt through both template compression and example absorption in Figure 3. During the initial fine-tuning phase, the input prompt will fully incorporate the template and 10-shot examples for the NL2F dataset. After a specified number of training iterations, the template will undergo compression at a rate of 0.3, and the number of examples will be reduced to five. This modified prompt is then used for the intermediate stage of fine-tuning. In the final phase, the template and few-shot examples are removed from the training prompt. It is important to note that the query remains unchanged throughout the entire progressive fine-tuning process. The prompt used in the last stage, which consists solely of the query, will also serve as the input for subsequent model inference. This method enables the fine-tuned language model to perform zero-shot inference without the need for an instruction or document template.
Appendix D Prompts
D.1 GPT-4 Generation (Baseline) instruction
For detailed prompts, please refer to Figure 4.
@@ D.1 GPT-4 Generation (Baseline) instruction
D.2 Prompts of PromptIntern on NL2Code
For detailed prompts of each dataset, please refer to Figures 5,6,7.